OpenStudy (anonymous):

can someone solve this problem for me ? find dy/dx (x,y)=(3,3) by implicit differentiation given: x^3+y^3=6xy

OpenStudy (sidsiddhartha):

just differentiate both sides using chain rule

OpenStudy (perl):

3x^2 + 3y^2 * y ' = 6* ( 1*y + x*y' )

OpenStudy (anonymous):

Im having trouble understanding implicit differentiation. I make it all the way to the point where I have to separate the dy/dx... I thought seeing someone else's work would help me see where I was going wrong... ;/

OpenStudy (sidsiddhartha):

the function is \[x^3+y^3=6xy\\differentiate~both~sides\\3x^2+3y^2*\frac{ dy }{ dx }=6x \frac{ dy }{ dx }+6y\]=

OpenStudy (sidsiddhartha):

ok? do u understand what i did

OpenStudy (anonymous):

The steps I do. i'm confused on why you kept the dy/dx on the 3y^2 and 6x... btw, thanks for helping me

OpenStudy (sidsiddhartha):

ok when you are differentiating y terms with respect to x, u have to include a dy/dx with it

OpenStudy (sidsiddhartha):

take a simple example \[f(x)=siny\\then\\f'(x)=cosy*\frac{ dy }{ dx}\]

OpenStudy (sidsiddhartha):

like this ok?

OpenStudy (anonymous):

Yeah! Thanks!

OpenStudy (sidsiddhartha):

can u do the rest? its just algebra ok!!

OpenStudy (sidsiddhartha):

take terms with dy/dx in any side then take common

OpenStudy (anonymous):

yes, i worked out the right answer :)

OpenStudy (sidsiddhartha):

well done :)

OpenStudy (perl):

@sidsiddhartha sid theres a mistake you wrote f(x) = sin y , it should be f(y) = sin y

OpenStudy (sidsiddhartha):

yup my bad it will be f(y)=siny as siny is a function of y thanks @perl

OpenStudy (perl):

given f(y(x)) (f(y(x)) ' = f ' (y(x)) * y' (x) or more simply [f(y)] ' = f ' (y) * dy/dx

OpenStudy (perl):

no problem :D